State, giving reasons, whether the following statements are True or False:
(1) An electron moves in a magnetic field of 1T with velocity 5.0 × 10^5
ms^–1
in the direction of
the magnetic field. The magnitude of the force experienced by the electron is 9.5 × 10^–14N.
1
Expert's answer
2020-02-17T09:20:52-0500

Solution. A moving electron in a magnetic field is affected by a force whose magnitude is equal to


"F=qvBsin\alpha"

where q=1.6 ×*10-19C is electron charge; v=5.0 × 10^5 ms^–1 is electron speed; B=1T is the magnetic field; α is the angle between the direction of electron velocity and the magnetic field. According to the conditions of the problem


"\alpha=0"

Therefore the magnitude of the force is equal to F=0 N. The statement is false.

Answer. False.